# Passive iFIR filters for data-driven control

### v1

@@ Abstract
I Introduction
@@
From simple prototyping to advanced control applications, a PID controller goes a long way due to its simple structure, intuitive tuning, and the presence of integral action for perfect regulation. In addition, with passive plants, PID control leads to inherently stable closed loops, even when the plant is uncertain and its model is approximated. This makes PID control one of the most common techniques in robotics and electronics [1, 2].
@@
However, in many control applications, a stabilizing feedback action must also satisfy complex performance requirements. These are often specified as desired input/output behaviors, in the form of a set of data or a desired transfer function. In this setting, PID control shows limitations [3, 4, 5]. Ideally, we want a more flexible controller that also retains the simplicity of PID control and its fundamental properties, namely, integral action and passivity.
@@
In this paper, we investigate the design of iFIR controllers, which combine the parallel action of a finite impulse response filter (FIR) and an integrator, for perfect regulation. Similar formulations, also based on cascade interconnections, can be found in [6, 7, 8, 9]. The idea is to replace the proportional and derivative action of the PID controller with the richer action of a FIR filter. We retain the integrator for perfect regulation. We also restrict the FIR filter to be passive, to guarantee wide applicability in the control of electro-mechanical systems and robotics [1, 2]. With integral action and passivity, our hypothesis is that iFIR controllers provide a more flexible alternative to PID control, when combined with data-driven optimal tuning.
@@
We propose a practical design approach based on model-free data-driven methods. Passivity is enforced through constrained optimization. The data-driven design of iFIR controllers is based on virtual reference feedback tuning (VRFT) [10], which reduces the matching problem with a target reference model to a least-squares fitting. No plant model is needed. The novelty of the paper is in endowing virtual reference feedback tuning with additional constraints for passivity. These are based on the Kalman-Yakubovich-Popov (KYP) lemma, on the Toeplitz operator of the iFIR controller, and on the positive realness of its transfer function. The result is a design approach based on convex constrained optimization, that can be solved with standard tools, such as CVX [11, 12, 13]. In what follows, we show that passivity constraints based on Toeplitz and positive realness formulations scale better with the size of the filter, strongly improving on the quadratic complexity of the KYP lemma.
@@
When combined with passive plants, passive iFIR control leads to a neat separation between closed-loop stability and closed-loop performance. Data scarcity and low-quality data do not affect the stability of the closed loop, which is structurally guaranteed via passivity. The quality and quantity of data still matter for performance, to achieve good matching with the desired reference model. This separation between stability and performance is novel. For instance, to guarantee stability, other data-driven approaches require persistency of excitation of the probing signal [3, 14]; or provide asymptotic guarantees in the limit of infinite data length [7, 5, 15]. The idea of enforcing passivity combined to data-driven approximation is borrowed by the general kernel approach to nonlinear modeling offered by [16]. There, passivity is achieved through regularization and scattering transform. We achieve it through constrained optimization, in the narrower setting of iFIR architectures.
@@
The paper is structured as follows. The iFIR control design based on VRFT is illustrated in Section 2. Passivity constraints are extensively discussed in Section 3. Section 4 provides an example of closed-loop control with a linear and nonlinear plants. The conclusions follow.
II Data-driven design of iFIR controllers
@@
$C(z)\ =\ \underbrace{\frac{\gamma T_{s}}{1-z^{-1}}}_{\mbox{integrator}}+\underbrace{\ \sum_{k=0}^{m-1}g_{k}z^{-k}}_{\mbox{FIR part}}$ (1)
where the integrator is discretized using backward Euler discretization with sampling period $T_{s}$. $\gamma$ is the integral gain. Indeed, $C$ shows a generalized PID structure, where the proportional and derivative terms are replaced by a FIR filter.
$C$ can be easily designed from data, following the approach of virtual reference feedback tuning [10]. From the block diagram in Figure 1, we represent the desired closed-loop performance with a time invariant (possibly nonlinear) reference model $M_{r}$. $P$ is a time-invariant single-input single-output plant. The plant is not necessarily linear but we assume passivity from the input $u$ to the output $y$. Under the action of the generic reference $r$, the ideal closed-loop behavior satisfies $y\simeq y^{*}$.
@@
where $\{y(t)\}_{t=0,...,N-1}$ is the measured open-loop plant output when the plant is excited with a user-defined signal $\{u(t)\}_{t=0,...,N-1}$. The output signal $y$ can be obtained through experiments, which makes the approach suitable to settings where a model of the plant is not fully available.
@@
Following [10], for linear plants and linear reference models, it can be shown that the minimizer of (2) also minimizes
@@
Remark 1
The optimization problems in (2) and (3) are equivalent to a least-squares problem and hence can be solved efficiently. To show this, define $e=M_{r}^{-1}(y)-y$. Following [17, Equation 7.3], the optimization problem (2) (or (3)) with the controller (1) reads
$\min_{g\in\mathbb{R}^{m}}\left\|u-Eg+\gamma E_{\mathrm{int}}\right\|_{2}^{2}$ (5)
where $g=\begin{bmatrix}g_{0}&\dots&g_{m-1}\\ \end{bmatrix}^{T}$, $u=\begin{bmatrix}u(0)&\dots&u(N-1)\\ \end{bmatrix}^{T}$,
@@
and $E_{\mathrm{int}}=T_{s}\begin{bmatrix}e(0)&e(0)+e(1)&\dots&\sum_{t=0}^{N-1}e(t)\end{bmatrix}^{T}.$
VRFT is an established and effective methodology. However, the data-driven optimal controller (2) does not guarantee stability of the closed-loop system. We tackle this problem via passivity.
III Passive iFIR controllers
For passive plants $P$, closed-loop stability is guaranteed if the iFIR controller is passive. Hence, we combine the least-square fitting (5) with specific constraints that enforce passivity. This leads to the derivation of the passive iFIR controller that best fit the desired input/output data (2).
The first formulation is a straightforward application of Kalman-Yakubovich-Popov (KYP) lemma.
Theorem 1 (KYP approach)
@@ III Passive iFIR controllers
$\begin{cases}\quad x(t+1)=Ax(t)+Bu(t)\\ \quad y(t)=Cx_{t}+Du(t)\end{cases}$ (6)
where $A\in\mathbb{R}^{m-1\times m-1},B\in\mathbb{R}^{m-1},C\in\mathbb{R}^{m-1},D\in\mathbb{R}$ are given by
$A$ $=\begin{bmatrix}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ \vdots&&&&&\\ 0&0&0&0&\dots&1\\ 0&0&0&\dots&&0\end{bmatrix},\quad B=\begin{bmatrix}0\\ 0\\ \vdots\\ 0\\ 1\end{bmatrix}$ (17)
$C$ $=\begin{bmatrix}g_{m-1}\quad g_{m-2}\quad\dots\quad g_{2}\quad g_{1}\end{bmatrix},\quad D=g_{0}\,.$ (19)
Define the following linear matrix inequalities
@@
$P+P^{T}$ $>0$ (20b)
$\quad\begin{bmatrix}P-A^{T}PA&\frac{1}{2}C^{T}-A^{T}PB\\ \frac{1}{2}C-B^{T}PA&D-B^{T}PB\end{bmatrix}$ $\geq 0\,.$ (20e)
Then, the optimal iFIR controller given by (2) and (1) is passive, thus guarantees closed-loop stability.
Proof:
(1) are the standard matrix inequalities for passivity of discrete LTI systems. These are linear in the unknowns $C$ and $P$. The passivity of the iFIR controller follows from the KYP lemma, recalling that the parallel interconnection of passive systems is passive. Closed-loop stability follows from the passivity theorem. ∎
The attracting feature of Theorem 1 is that the constrained minimization problem can be handled efficiently by convex optimization solvers such as CVX, [11, 12, 13]. The main issue of (1) is that the number of unknowns scales as $o(m^{2})$. This may cause long computation times and numerical precision issues, leading to convergence issues when the order of the FIR part is high. The approach proposed below, overcomes these limitations.
@@
The first $n$ samples of the response of the FIR filter given by $g$ can be computed as the product between $\phi_{n}(g)$ and the (vectorized) input $e=[e(0),\dots,e(n-1)]^{T}$. It is thus not surprising that the passivity of the FIR filter is related with the positive semi-definiteness of the Toeplitz matrix $\phi_{n}(g)+\phi_{n}(g)^{T}$, as stated in the following theorem. In what follows the symbol $\underline{\sigma}$ denotes the minimal singular value of a matrix.
Theorem 2 (Infinite Toeplitz approach)
@@
Proof:
Like the proof of Theorem 1, we just need to show that the FIR part of the filter is passive. Consider the vectorized generic input $w_{n}=[w_{0},w_{1},\dots,w_{n}]^{T}$ and associated output $z_{n}=[z_{0},z_{1},\dots,z_{n}]^{T}$ of the FIR part, where $n\geq m$. For all $n\geq m$, we have
@@
Therefore,
$z_{n}^{T}w_{n}\geq 0\quad\equiv\quad\phi_{n}(g)^{T}+\phi_{n}(g)\geq 0.$
Thus, for $n\to\infty$, passivity requires (22b). ∎
The main issue of Theorem 2 is that (2) is not computationally tractable. This motivates the following relaxation.
Theorem 3 (Finite Toeplitz approach)
Given the iFIR (1) of order $m$, let $g=\{g_{k}\}_{k=0,...,m-1}$ be the finite coefficients of its FIR part. For any $\epsilon>0$, $\rho_{0}>0$, $0<\rho\leq 1$, and there exists $n^{*}\geq m$ such that for all $n\geq n^{*}$, the iFIR controller given by (2) and
@@
Proof:
Again, we just need to show the FIR part of the iFIR controller is passive.
For $n\to\infty$, the matrix $\phi_{n}(g)^{T}+\phi_{n}(g)$ is a $(m-1)-banded$ symmetric infinite dimensional Toeplitz matrix, [19, Theorem 2.1], [18, pp.37]. Hence, as shown in [18, Equation 4.5] and [19, p. 322], there exists an associated generating function $f$ given by
@@
where $\omega\in[-\frac{1}{2},\frac{1}{2})$. $\{g_{k}\}_{k=-\infty,...,\infty}$ is absolutely summable. Therefore, we can apply [18, Corollary 4.2] to obtain
@@
where $\operatorname*{ess\,inf}$ refers to essential infimum defined in [18, pp.40]. Since $|f(\omega)|\leq|g_{0}|+\sum_{k=-m+1}^{m-1}|g_{k}|$, it follows that the limit is bounded.
@@
for all $n\geq n^{*}$. Thus, for any given $n\geq n^{*}$, (23d) guarantees $\lim_{k\to\infty}\underline{\sigma}\!\left(\phi_{k}^{T}(g)+\phi_{k}(g)\right)\geq 0$. This implies that the FIR part of the iFIR controller is passive. ∎
$\rho_{0}$, and $\rho$ are not strictly needed for the proof of the theorem. However, these user-defined parameters determine the decay rate of the FIR coefficients, and can be used to reduce the Gibbs phenomenon due to sharp truncation of the impulse response of the FIR part of the controller. (23d) is a relaxation of (22b) based on a finite constraint plus additional bound $\epsilon$. The latter reduces as $n$ increases (and vice versa). Numerical tests show that, for large $n$, the Nyquist diagram of a FIR filter that satisfies $\phi_{n}(g)^{T}+\phi_{n}(g)\geq 0$ has a large portion of its locus on the right-half of the complex plane. To have the locus completely within the right-half plane, we can ‘shift’ the whole Nyquist diagram towards right using strictly input passivity, [20]. This is equivalent to take $\phi_{n}(g)^{T}+\phi_{n}(g)\geq\epsilon I$.
Theorem 3 provides a computationally tractable optimization problem. A passive iFIR can be obtained by solving the optimization iteratively, for increasing values of $n$: (i) we select $\rho_{0}$, $\rho$ and $\epsilon$ and solve the optimization problem for a given $n\geq m$; (ii) we test passivity of the iFIR (from the Nyquist plot or using any other test); (iii) in case of shortage of passivity, we select a larger $n$ (and possibly a larger $\epsilon$) and iterate.
Compared with the KYP approach of Theorem 1, the finite Toeplitz approach of Theorem 2 leads to an optimization problem with $o(m)$ unknowns, linear in the order $m$ of the iFIR filter. However, $\epsilon$ is not known and needs to be tuned iteratively. The following theorem overcomes this limitation. Given the coefficients $g={g_{k}}$, in what follows we use $G(e^{j\theta})$ to denote its frequency response $\sum_{k=0}^{m-1}g_{k}e^{-jk\theta}$.
Theorem 4 (Positive realness approach)
@@
Thus, from (30), it is easy to see that (27d) and (27e) guarantee $G(e^{j\theta})+G(e^{-j\theta})=f(\theta)\geq 0$. ∎
@@
Theorem 4 follows closely [21, Section III.B] but provides the exact bound (27e), to unlock a non-iterative design of the iFIR controller. The theorem illustrates that passivity can be enforced by constraining the sampled frequency response of the FIR to be larger than a certain constant $\epsilon$, which depends on the decay rate of the impulse response, the order of FIR part, and the number of sampling points in the frequency domain. This is equivalent to verify that the sampled Nyquist diagram of the FIR subcontroller has real part greater than $\epsilon$. In this case, $\epsilon$ bounds the maximal variation of the real part of the Nyquist diagram between each two samples.
The bound (27e) is small if the ratio between the size of the FIR part, $m$, and the number of sampling points, $M$, is small. Indeed, for large orders $m$, this implies that either the bound $\epsilon$ is conservative or the optimization problem has a large number of constraints, $M$. The trade-off between the computational complexity and conservativeness can be mitigated by replacing the exact bound (27e) with an heuristic search over $\epsilon$, based on an iteration similar to the one derived from Theorem 3.
Compared with Theorem 3, Theorem 4 leads to a one-shot design approach. In terms of complexity, Theorem 4 is more efficient than Theorems 1 and 3 since no LMIs are involved. (27d) can be implemented with $M+1$ linear inequality constraints of the form (28), for $\theta=\frac{q\pi}{M}$ and $q\in\{0,\dots,M\}$.
@@
Figure 2: Nyquist diagrams of Example IV-A. blue - target filter; black - identified passive FIR filter. Top: $C_{1}$ (left), $C_{1}$ and excess of passivity (right). Bottom: $C_{2}$ (left), $C_{3}$ (right).
IV Examples
IV-A Fitting of passive FIR filters
In this section we identify passive FIR filters to illustrate the effectiveness of (1), (3) and (4). Three target systems are considered:
$C_{1}(s)=\frac{1}{0.5s+1}\quad C_{2}(s)=\frac{1}{(0.5s+1)^{2}}\quad C_{3}(s)=\frac{1}{(0.5s+1)^{3}}.$ (31)
$C_{1}$ is passive. $C_{2}$ and $C_{3}$ show increase shortage of passivity. For each target, we take $\gamma=0$ and we use (5) paired to passivity constraints. The input/output data $(e,u)$ are generated as step response $u=C_{i}e$ for $i\in\{1,2,3\}$. Signals are low-pass filtered through $\frac{1}{0.2s+1}$ to improve the fitting.

### v2

@@ Abstract
{IEEEkeywords}
Data driven control, Identification for control, PID control, LMIs.
1 Introduction
[…]
@@ 1 Introduction
From simple prototyping to advanced control applications, a PID controller goes a long way due to its simple structure, intuitive tuning, and the presence of integral action for perfect regulation. In addition, with passive plants, PID control leads to inherently stable closed loops, even when the plant is uncertain and its model is approximated. This makes PID control one of the most common control techniques in industry [1, 2, 3].
@@
In many control applications, a stabilizing feedback action must also satisfy complex performance requirements. These are often specified as desired input/output behaviors, in the form of a set of data or a desired transfer function. In this setting, the PID control shows limitations [4, 5]. Ideally, we want a controller that retains the simplicity of PID control and its fundamental properties, namely integral action and passivity, but is more flexible.
@@
In this paper, we investigate the design of iFIR controllers, which combine the parallel action of a finite impulse response filter (FIR) and an integrator for perfect regulation. Similar formulations, also based on cascade interconnections, can be found in [6, 7, 8, 9]. The idea is to replace the proportional and derivative action of the PID controller with the richer action of a FIR filter. We retain the integrator for perfect regulation. We also restrict the FIR filter to be passive, to guarantee wide applicability in the control of electro-mechanical systems and robotics [1, 2, 3]. With these features, our hypothesis is that iFIR controllers provide a more flexible alternative to PID control, when combined with data-driven optimal tuning.
@@
We propose a model-free data-driven design approach based on virtual reference feedback tuning (VRFT) [10], which reduces the matching problem with a target reference model to a least-squares fitting. No plant model is needed. Passivity is enforced through constrained optimization. The novelty of the paper is in endowing virtual reference feedback tuning with additional constraints for passivity. These are based on the Kalman-Yakubovich-Popov (KYP) lemma, on the Toeplitz operator of the iFIR controller, and on the positive realness of its transfer function. The result is a design approach based on convex constrained optimization, that can be solved with standard tools, such as CVX [11, 12, 13]. In what follows, we show that passivity constraints based on Toeplitz and positive realness formulations scale better with the size of the filter, strongly improving the complexity of the KYP lemma.
@@
For passive plants, passive iFIR control design separates closed-loop stability and closed-loop performance. Data scarcity and low-quality data do not affect the stability of the closed loop, which is structurally guaranteed via passivity. The quality and quantity of data matter for the performance, to achieve good matching with the desired reference model. This separation between stability and performance is novel. To guarantee stability, other data-driven approaches require the probing signal to be persistently exciting [4] or informative [14], or provide asymptotic guarantees in the limit of infinite data length [7, 5, 15].
The idea of enforcing passivity combined with a data-driven approximation is inspired by the kernel approach to nonlinear modeling offered by [16]. There, passivity is achieved through regularization and the scattering transform. We achieve it through constrained optimization, in the narrower setting of iFIR architectures. Passive filter design has been also explored in [17] and [18]. These approaches show strong similarities with our formulation. Our contribution is in the combination with VRFT for control design purposes (Theorems 3.1, 3.5, and 3.7), in the formulation based on Toeplitz matrices (Theorems 3.3 and 3.5), and in the characterization of exact bounds on (sampled) frequency-based constraints (Theorem 3.7), which lead to a non-iterative design.
@@
The paper is structured as follows. The iFIR control design based on VRFT is illustrated in Section 2. Passivity constraints are extensively discussed in Section 3. Section 4 provides an example of closed-loop control with linear and nonlinear plants. The conclusions follow.
2 Data-driven design of iFIR controllers
@@
$C(z)\ =\ \underbrace{\frac{\gamma T_{s}}{1-z^{-1}}}_{\mbox{integrator}}+\underbrace{\ \sum_{k=0}^{m-1}g_{k}z^{-k}}_{\mbox{FIR filter}}$ (1)
where the integrator is discretized using backward Euler discretization with sampling period $T_{s}$. $\gamma$ is the integral gain. In fact, $C$ shows a generalized PID structure, where the proportional and derivative terms are replaced by a FIR filter.
$C$ can be easily designed from data, following the approach of virtual reference feedback tuning [10]. In Figure 1, we represent the desired closed-loop performance with a time-invariant (possibly nonlinear) reference model $M_{r}$. $P$ is a time-invariant single-input single-output plant. The plant is not necessarily linear but we assume passivity from the input $u$ to the output $y$ (e.g. Euler-Lagrangian systems and port-Hamiltonian systems [1, 19]). For any given reference $r$, the ideal closed-loop behavior satisfies $y\simeq y^{*}$.
@@
where $\{y(t)\}_{t=0,...,N-1}$ is the measured open-loop plant output when the plant is excited with a user-defined signal $\{u(t)\}_{t=0,...,N-1}$. The output signal $y$ can be obtained by experiments, making the approach suitable for settings where a plant model is not fully available.
@@
Following [10], for linear plants and linear reference models, the minimizer of (2) also minimizes
@@
The optimization problems in (2) and (3) are equivalent to a least-squares problem and hence can be solved efficiently. For instance, take $e=M_{r}^{-1}(y)-y$. From [20, Eq. 7.3], the optimization problem (2) (or (3)) with the controller (1) reads
$\min_{g\in\mathbb{R}^{m}}\left\|u-Eg{\color[rgb]{0,0,0}-}\gamma E_{\mathrm{int}}\right\|_{2}^{2}$ (5)
where $\small{g=\begin{bmatrix}g_{0}&\dots&g_{m-1}\\ \end{bmatrix}^{T}}$, $\small{u=\begin{bmatrix}u(0)&\dots&u(N-1)\\ \end{bmatrix}^{T}}$,
@@
and $\small{E_{\mathrm{int}}=T_{s}\begin{bmatrix}e(0)&e(0)+e(1)&\dots&\sum_{t=0}^{N-1}e(t)\end{bmatrix}^{T}.}$ If the input is persistently exciting, then $E$ will be full rank and a unique solution exists [4, Section 3.2]. In what follows, in our extension of VRFT based on passive iFIR controller, uniqueness of solution is not necessary to guarantee closed-loop stability. Therefore, this rank condition will not be considered again.
3 Passive iFIR controllers
Although VRFT is an established and effective design methodology, the data-driven optimal controller (2) does not guarantee closed-loop stability. For passive plants $P$, we tackle this issue through passivity. We combine the least-square fitting (5) with specific constraints that enforce passivity. This leads to the derivation of the passive iFIR controller that best fits the desired input / output data (2).
The first formulation is a straightforward application of the Kalman-Yakubovich-Popov (KYP) lemma.
Theorem 3.1 (KYP approach)
@@ 3 Passive iFIR controllers
$\begin{cases}\quad x(t+1)={\color[rgb]{0,0,0}A_{c}}x(t)+{\color[rgb]{0,0,0}B_{c}}{\color[rgb]{0,0,0}w(t)}\\ \quad{\color[rgb]{0,0,0}z(t)}={\color[rgb]{0,0,0}C_{c}}{\color[rgb]{0,0,0}x(t)}+{\color[rgb]{0,0,0}D_{c}}{\color[rgb]{0,0,0}w(t)}\end{cases}$ (6)
where $w(t),z(t)$ are the input and output of the FIR part. $A_{c}\in\mathbb{R}^{m-1\times m-1},B_{c}\in\mathbb{R}^{m-1},C_{c}\in\mathbb{R}^{m-1},D_{c}\in\mathbb{R}$ are given by
${\color[rgb]{0,0,0}A_{c}}$ $=\begin{bmatrix}0&1&0&0&\dots&0\\ 0&0&1&0&\dots&0\\ \vdots&&&&&\\ 0&0&0&0&\dots&1\\ 0&0&0&\dots&&0\end{bmatrix},\quad{\color[rgb]{0,0,0}B_{c}}=\begin{bmatrix}0\\ 0\\ \vdots\\ 0\\ 1\end{bmatrix}$
${\color[rgb]{0,0,0}C_{c}}$ $=\begin{bmatrix}g_{m-1}\quad g_{m-2}\quad\dots\quad g_{2}\quad g_{1}\end{bmatrix},\quad{\color[rgb]{0,0,0}D_{c}}=g_{0}\,.$
The optimal iFIR controller given by (2) constrained to the following LMIs in the unknown $X\in\mathbb{R}^{m-1\times m-1}$, $C_{c}$, and $D_{c}$
@@
$X=X^{T}$ $>0$ (7b)
$\quad\begin{bmatrix}X-A_{c}^{T}XA_{c}&C_{c}^{T}-A_{c}^{T}XB_{c}\\ C_{c}-B_{c}^{T}XA_{c}&D_{c}+D_{c}^{T}-B_{c}^{T}XB_{c}\end{bmatrix}$ $\geq 0\,$ (7c)
$D_{c}+D_{c}^{T}$ $\geq 0.$ (7d)
is passive, thus guarantees closed-loop stability.
Proof 3.2.
(7) are the standard matrix inequalities for passivity of discrete LTI systems. These are linear in the unknowns $C_{c}$ and $X$. The passivity of the iFIR controller follows from the KYP lemma [18, Thm. 1], recalling that the parallel interconnection of passive systems is passive. Closed-loop stability follows from the passivity theorem.
The attractive feature of Theorem 3.1 is that the constrained minimization problem can be handled efficiently by convex optimization solvers such as CVX, [11, 12, 13]. The main issue of (7) is that the number of unknowns scales as $o(m^{2})$. This may cause long computation times and numerical precision issues, leading to convergence issues when the order of the FIR part is high. The approach proposed below, overcomes these limitations.
@@
The first $n$ samples of the response of the FIR filter given by $g$ can be computed as the product between $\phi_{n}(g)$ and the (vectorized) input $w_{n}=\begin{bmatrix}w(0)&w(1)&\dots&w(n-1)\end{bmatrix}^{T}.$ It is thus not surprising that the passivity of the FIR filter is related to the positive semi-definiteness of the Toeplitz matrix $\phi_{n}(g)+\phi_{n}(g)^{T}$, as stated in the following theorem. In what follows, the symbol $\underline{\sigma}$ denotes the minimal singular value of a matrix.
Theorem 3.3 (Infinite Toeplitz approach).
@@
Proof 3.4.
Like the proof of Theorem 3.1, we just need to show that the FIR part of the filter is passive. Consider any vectorized input $w_{n}=\begin{bmatrix}w(0)&w(1)&\dots&w(n-1)\end{bmatrix}^{T}$ and associated output $z_{n}=\begin{bmatrix}z(0)&z(1)&\dots&z(n-1)\end{bmatrix}^{T}$ of the FIR part, where $n\geq m$. Denote by $w_{\infty},z_{\infty}$ the related (infinite length) vectors when $n\to\infty$. For passivity, we require $z_{\infty}^{T}w_{\infty}\geq 0$ for all $w_{\infty}$. For all $n\geq m$, we have
@@
Therefore, saying that $z_{n}^{T}w_{n}\geq 0$ for all inputs $w_{n}$ (and related outputs $z_{n}$) is equivalent to the matrix inequality
$\phi_{n}(g)^{T}+\phi_{n}(g)\geq 0.$
Thus, for $n\to\infty$, (9b) is equivalent to $z_{\infty}^{T}w_{\infty}\geq 0$.
The main issue of Theorem 3.3 is that (9) is not computationally tractable. This motivates the following relaxation.
Theorem 3.5 (Finite Toeplitz approach).
Given the iFIR (1) of order $m$, let $g=\{g_{k}\}_{k=0,...,m-1}$ be the finite coefficients of its FIR part. For any $\epsilon>0$, $\rho_{0}>0$, $0<\rho\leq 1$, there exists $n^{*}\geq m$ such that, for all $n\geq n^{*}$, the iFIR controller given by (2) and
@@
Proof 3.6.
We show that the FIR part of the controller is passive. For $n\to\infty$, the matrix $\phi_{n}(g)^{T}+\phi_{n}(g)$ is a $(m-1)-banded$ symmetric infinite dimensional Toeplitz matrix, [22, Thm. 2.1], [21, p. 37]. Hence, as shown in [21, Eq. 4.5] and [22, p. 322], there exists an associated generating function $f$ given by
@@
where $\omega\in[-\frac{1}{2},\frac{1}{2})$. $\{g_{k}\}_{k=-\infty,...,\infty}$ is absolutely summable. Therefore, we can apply [21, Corollary 4.2] to obtain
@@
where $\operatorname*{ess\,inf}$ refers to essential infimum defined in [21, p. 40]. Since $|f(\omega)|\leq|g_{0}|+\sum_{k=-m+1}^{m-1}|g_{k}|$, it follows that the limit is bounded.
@@
for all $n\geq n^{*}$. Thus, for any given $n\geq n^{*}$, (10d) guarantees $\lim_{k\to\infty}\underline{\sigma}\!\left(\phi_{k}^{T}(g)+\phi_{k}(g)\right)\geq 0$. This implies that the FIR part of the iFIR controller is passive.
$\rho_{0}$, and $\rho$ are not strictly needed for the proof of the theorem. However, these user-defined parameters determine the decay rate of the FIR coefficients, and can be used to reduce the Gibbs phenomenon due to sharp truncation of the impulse response of the FIR part of the controller. (10d) is a relaxation of (9b) based on a finite constraint plus an additional bound $\epsilon$. The latter reduces as $n$ increases (and vice versa). Numerical tests show that, for large $n$, the Nyquist diagram of a FIR filter that satisfies $\phi_{n}(g)^{T}+\phi_{n}(g)\geq 0$ has a large portion of its locus on the right-half of the complex plane. To have the locus completely within the right-half plane, we can ‘shift’ the whole Nyquist diagram right using strictly input passivity, [23]. This is equivalent to taking $\phi_{n}(g)^{T}+\phi_{n}(g)\geq\epsilon I$.
Theorem 3.5 provides a computationally tractable optimization problem. A passive iFIR can be obtained by solving the optimization iteratively, for increasing values of $n$: (i) we select $\rho_{0}$, $\rho$ and $\epsilon$ and solve the optimization problem for a given $n\geq m$; (ii) we test passivity of the iFIR (from the Nyquist plot or using any other test); (iii) in case of shortage of passivity, we select a larger $n$ (and possibly a larger $\epsilon$) and iterate.
Compared with the KYP approach of Theorem 3.1, the finite Toeplitz approach of Theorem 3.3 leads to an optimization problem with $o(m)$ unknowns, linear in the order $m$ of the iFIR filter. However, for a fixed $\epsilon$ (or $n$), the value of $n$ (or $\epsilon$) that leads to passivity is bounded by some finite but unknown $n^{*}$ (by some finite $\epsilon^{*}$), thus needs to be tuned iteratively. The following theorem overcomes this limitation. Given the coefficients $g={g_{k}}$, in what follows we use $G(e^{j\theta})$ to denote its frequency response $\sum_{k=0}^{m-1}g_{k}e^{-jk\theta}$.
Theorem 3.7 (Positive realness approach).
@@
Thus, from (17), it is easy to see that (14d) and (14e) guarantee $G(e^{j\theta})+G(e^{-j\theta})=f(\theta)\geq 0$.
@@
Theorem 3.7 follows closely [17, Section III.B] and [18, Section 2] but guarantees passivity with a finite length of training data and a finite number of constraints in the context of data-driven control design. Compared with Theorem 3.5, Theorem 3.7 leads to a one-shot design approach. The theorem illustrates that passivity can be enforced by constraining the sampled frequency response of the FIR to be larger than a certain constant $\epsilon$, which depends on the decay rate of the impulse response, the order of the FIR part, and the number of sampling points in the frequency domain. This is equivalent to verifying that the sampled Nyquist diagram of the FIR subcontroller has a real part greater than $\epsilon$. In this case, $\epsilon$ bounds the maximal variation of the real part of the Nyquist diagram between each two samples.
From the proof of Theorem 3.7, it can be seen that the bound (14e) on $\epsilon$ is conservative. Conservativeness is reduced by increasing the number of samples $M$, that is, by enforcing a larger number of constraints. The trade-off between the computational complexity and conservativeness can be mitigated by replacing the bound (14e) with an heuristic search over $\epsilon$, based on an iteration similar to the one derived from Theorem 3.5. In terms of complexity, Theorem 3.7 is more efficient than Theorems 3.1 and 3.5 since there are no LMIs involved. Furthermore, (14d) can be implemented with $M+1$ linear inequality constraints of the form (15), for $\theta=\frac{q\pi}{M}$ and $q\in\{0,\dots,M\}$.
Remark 3.9.
The constrained optimization (2) subject to passivity constraints is always feasible. For instance, for any $\epsilon^{*}\in\mathbb{R}$ and $\gamma\geq 0$, consider the iFIR controller in (1) given by $g_{0}\geq\epsilon^{*}$ and $g_{k}=0,\forall k>0$. If $\epsilon^{*}\geq 0$, this iFIR controller is compatible with (7). If $\epsilon^{*}$ is larger than the $\epsilon$ chosen in (10) or (14e), then this iFIR controller is compatible with (10) or (14), respectively.
Remark 3.10.
The phase correction that a passive linear controller can introduce is limited to ±$\pm 90$ degrees. This poses a limit on the achievable performance. The level of conservativeness depends on the performance requirement, based on the selected reference model $M_{r}$.
@@
Figure 2: Nyquist diagrams of Example 4.1. Blue - target filter; black - identified passive FIR filter. Top: $C_{1}$ (left), $C_{1}$ and excess of passivity (right). Bottom: $C_{2}$ (left), $C_{3}$ (right).
4 Examples
4.1 Fitting of passive FIR filters
We identify passive FIR filters to illustrate the effectiveness of (7), (10) and (14) in dealing with the targets
$C_{q}(s)=\frac{1}{(0.5s+1)^{q}}\qquad q\in\{1,2,3\},$ (18)